OpenStudy (anonymous):

Which is an equation in slope-intercept form for the line passing through (–1, 5) and (5, 8).

OpenStudy (anonymous):

@wolf1728

OpenStudy (anonymous):

Can you help?

OpenStudy (anonymous):

Yes. So first we need to find the slope. To do that, we do 5-8/-1-5. Can you simplify that?

OpenStudy (anonymous):

so -3/-6

OpenStudy (anonymous):

thats -1/2

OpenStudy (anonymous):

Good, but remember that both negative signs cancel out. So it will be only 1/2. NExt we put in point slope form using the first point, y-5=1/2(x-8) Now distribute the right side. What do you get?

OpenStudy (anonymous):

Wait how do I do that?

OpenStudy (anonymous):

I mean how do I distribute the right side?

OpenStudy (anonymous):

To distribute the right side, do1/2 times x, and 1/2 times -8.

OpenStudy (anonymous):

so 1/2x and -4?

OpenStudy (anonymous):

Yes. :) So now we have y-5=1/2x-4. Now add 5 to both sides. The 5 on the left side will cancel out, so what do you have now?

OpenStudy (anonymous):

y=1/2x+1?

OpenStudy (anonymous):

That's your equation. :D

OpenStudy (anonymous):

A. y=-1/2x+11/2 B. y=1/2x-11/2 C. y=1/2x+11/2 D. y=-1/2x-11/2 I have to choose from these....

OpenStudy (anonymous):

Maybe its C.?

OpenStudy (anonymous):

LOL I see what I did wrong. Sorry. The equation should have been y-5=1/2(x+1) Lets use the point (-1,5) y-5=1/2x+0.5 (distribute) y=1/2x+5.5 (add 5 to both sides)' or y=1/2x+11/2 Sorry for the previous mistake

OpenStudy (anonymous):

Its ok thank you so much do you think I can tag you in another question for more help? :) You explain things so well and walk me through all the steps!!

OpenStudy (anonymous):

Sure. :D

OpenStudy (anonymous):

Ok let me close this one and tag you so look out for it! :)